### v1

@@ 6   Factors underlying effective GEM-V stories and voxels
Figure 4: Analyzing factors that impact explanation-mediated validation. To evaluate whether GEM-V succeeds in generating effective stimulus stories, we assessed the driving paragraphs of the stimulus. (a) To confirm that generated paragraphs match the explanation used to construct them, a matching score was computed for each explanation and paragraph by using an LLM to evaluate the fraction of trigrams in the paragraph that are relevant to the paragraph’s generating explanation and then z-scoring the result. Each driving paragraph showed a strong match with its generating explanation. Plot shows one subject (S02), similar plots for other subjects are shown in Section A.4. (b) To confirm that each driving paragraph effectively drives its corresponding encoding model predictive performance, we computed the predicted response in each selected voxel to each generated paragraph. This revealed strong matches for most voxels, along with some matches between driving paragraphs and voxels with semantically similar explanations, e.g. directions and locations. Plot shows one subject (S02), similar plots for other subjects are shown in Section A.4. (c) After running the fMRI driving experiment, we found that a key factor determining whether a voxel is driven well by the GEM-V framework was the stability score for the voxel, i.e. the correlation between the n-gram rankings provided by the LLaMA-based encoding model and the OPT-based encoding model. (d) Another important factor for eliciting increased driving responses is the presence of key n-grams in the driving paragraphs. These n-grams induce a standard hemodynamic response curve that peaks at around 6 seconds, yielding a significant increase ($p=0.009$; one-sided t-test). (e) Finally, to test whether the driving results were sensitive to the particular voxels that were selected, we evaluated whether the GEM-V stories drove alternative voxels in each subject that were assigned the same explanation as the target voxels being driven. Both the targed voxels and alternative voxels also showed significantly increased driving responses ($p<0.05$; permutation test, FDR-corrected).
@@
The experiments thus far showed that GEM-V can effectively explain and drive a variety of locations throughout cortex, but fails to successfully drive some voxels. To understand these limitations, we evaluated factors that influence driving performance, beginning with factors surrounding the generation of story stimuli.
@@
Given that driving failures are neither the result of stimulus-explanation mismatches nor stimulus-model mismatches, we concluded that they must largely stem from limitations in the encoding models. We only applied GEM-V to voxels with high encoding model test performance, but this does not guarantee that the encoding model is accurate. Model performance is measured on a dataset that is limited in scope and that comprises real, noisy data. Both of these factors could cause a voxel to appear well-predicted despite having an inaccurate encoding model, or to appear poorly-predicted despite having an accurate encoding model. Since these limitations cannot be addressed directly, we investigated factors that improve confidence in the generated explanations by assessing their stability, a key principle underlying effective statistical interpretation according to the predictability, computability, and stability (PCS) framework [14, 29].
@@ 7   Discussion
This work is related to works that have generated stimuli to drive neurons in the visual cortex [38, 15, 39], as well as more recent work that has used decoding to provide insight into visual cortex selectivity [40]. In language fMRI, the closest work is Tuckute et al. 2023, which demonstrated the ability to drive the language network as a whole (rather than for individual voxels or ROIs) using non-generative sampling techniques as demonstrated in 13. In comparison to these prior works, ours is the first to demonstrate driving at an explanatory level, not only identifying stimuli that can drive a particular voxel or ROI, but also hypothesizing the shared semantic features that cause the underlying activity. The methodology here runs parallel to methods for causal interventions in mechanistic interpretability [41, 42, 43], which have been applied to understanding artificial neural networks.
@@
Our work can be applied to generating a new class of stimuli for online, causal experiments of language processing. We consider this work a piece of a larger online feedback system that continually refines its understanding of brain selectivity by generating hypotheses, testing these hypotheses with new experiments, and refining the hypotheses based on the experimental results. As an example, such a framework could be an especially effective tool in lesion case studies where general observations must be made efficiently from relatively little data.
@@
While effective, GEM-V has several limitations. Most notably, it focuses on a single explanation, missing voxel activity that is polysemantic, i.e. that is not driven well by any single explanation. See Section A.2 for a more detailed discussion of this limitation and how future work might overcome it. Second, the effectiveness of GEM-V depends on staying within the manifold of stimuli trained on during encoding model training. This limitation is explored further in A.6 where we show that poorly generated, off-manifold stimuli are more difficult to predict and yield inferior driving outcomes. GEM-V is also sensitive to the properties of the underlying encoding models–using different encoding models could yield different explanations which might be equally valid. Thus we must be cautious not to interpret the explanations given by GEM-V as uniquely capturing the functional properties of a voxel or ROI. However, this issue will become less severe over time as better and more diverse encoding models become available.
@@
The work here opens the door to a great deal of future research. One avenue should explore whether the degree of granularity of the explanation that GEM-V generates is appropriate for cortical language selectivity. For example, one of the explanations proposed by GEM-V is birthdays, but it would be somewhat surprising to have a specifically birthday-selective area in human cortex. Rather, it is more plausible that such a voxel is selective for something more generic, such as social gatherings [44]. Such ambiguities are of the type that are liable to be resolved by higher-fidelity modelling or iterative causal experiments, such as the hypothesis-refinement feedback system proposed above. GEM-V is also fundamentally limited by the spatiotemporal granularity of the scanning modality we test it on, fMRI. Applying GEM-V with faster or more fine-grained scanning modalities, such as 7T laminar fMRI, ECoG, or NeuroPixels, could reveal selectivity patterns that cannot be effectively observed through fMRI.
@@
GEM-V demonstrates a general framework for aligning the data-driven predictions of models to the theory-oriented hypotheses of classical science. In this paper, we show that applying techniques from modern LLM interpretability research can improve our understanding of cognitive brain processes. Through simple prompting-based methods, we successfully recover semantic selectivity hypotheses that have previously been the providence of painstaking controlled experiments created over decades, and open the door to many more similar results. Our framework is extensible to other modalities such as vision and - requiring only naturalistic data from a single subject - can be easily adapted for use in case study analyses. Furthermore, GEM-V is data-driven and requires minimal human input, countering a frequent criticism of such theories as cherry-picked or driven by preexisting scientific biases. GEM-V will help move toward a future with increasingly nuanced, operationalized, and falsifiable explanations in language neuroscience.
@@ 8   Methods
Framework for generating explanations GEM-V yields a short, natural-language explanation describing what elicits the strongest response from each of the 1,500 selected voxel encoding models. To obtain this explanation, we follow the summarize and score (SASC) framework introduced in previous work [17] (see Fig. 1b). SASC first generates candidate explanations (using GPT-4 [16]) based on the n-grams ($n=1,2,3$) that elicit the most positive response from the LLaMA encoding model. Each candidate explanation is then evaluated by generating synthetic data based on the explanation and testing the response of the encoding model to the data.
@@
Before running follow-up experiments, we checked that each paragraph’s text matched its generated explanation (Fig. 4a). This match was measured by prompting an LLM [47] to evaluate the fraction of trigrams in the paragraph that are relevant to the paragraph’s generating explanation. We also validated that each story paragraph drives the encoding model for its corresponding voxel (Fig. 4b). We generated 8 different stories by changing the random seed for each subject and kept the best 2 for S01, the best 6 for S02 (the pilot subject), and the best 2 for S03.
@@ A.1 Prompting details
OPA Direction and location descriptions Avoid mentioning any specific location names (like ”New York” or ”Europe”).
OPA-only Self-reflection and growth Avoid mentioning any specific location names (like ”New York” or ”Europe”).
PPA Scenes and settings Avoid mentioning any specific location names (like ”New York” or ”Europe”).
PPA-only Unappetizing foods Avoid mentioning any specific location names (like ”New York” or ”Europe”).
@@ A.2 Driving polysemantic voxels
We additionally run experiments seeking to drive multiple descriptions corresponding to the same voxel. These experiments generate and test GEM-V stories in the same manner as testing individual explanations (Fig. 1c) but differ in the way they source voxel descriptions. In our first investigation (with subject S02), we select two explanations that both come from SASC, in the case that its summaries of the top ngrams yield two distinct explanations. In our second investigation (now with subject S03), we instead select two explanations that each come from a different encoding model. In both cases, each story tests 8 voxels with two explanations each. We average the results over two stories per subject.
@@ A.3 Story-level breakdowns
In this section, we show results for GEM-V at the level of individual stories. Fig. 6 shows the mean driving voxel response across all subjects and settings. Fig. 7 stratifies these synthetic stories based on the underlying prompt used to generate the stories and Fig. 8 shows results for these stories at driving voxel clusters rather than individual voxels.
@@
Table 2: Head motion for different subjects during GEM-V stories. S01 shows substantially larger motion than the other two subjects.
@@ A.5 Selectively driving pairs of voxels
Useful and explanatory scientific theories are robust to unrelated distributional shifts. For example, if a region or voxel is selective for food concepts, then that region should remain selective for food concepts even if those concepts cooccur with unrelated semantic concepts. In order to test this independence of our generated explanations, we test whether GEM-V can be used to simultaneously drive pairs of voxels in different areas across the brain.
@@ A.6 Checkerboard driving
A.7 Semantic explanation average constrast maps [^1] Full set of flatmaps can also be found at github.com/microsoft/automated-explanations

### v2

@@ 7   Factors affecting GCT’s ability to drive a voxel
Figure 4: Analyzing factors that impact explanation-mediated validation. To evaluate whether GCT succeeds in generating effective stimulus stories, we assessed the driving paragraphs of the stimulus. (a) To confirm that generated paragraphs match the explanation used to construct them, a matching score was computed for each explanation and paragraph by using an LLM to evaluate the fraction of trigrams in the paragraph that are relevant to the paragraph’s generating explanation and then z-scoring the result. Each driving paragraph showed a strong match with its generating explanation. Plot shows one subject (S02), similar plots for other subjects are shown in Section A.4. (b) To confirm that each driving paragraph effectively drives its corresponding encoding model predictive performance, we computed the predicted response in each selected voxel to each generated paragraph. This revealed strong matches for most voxels, along with some matches between driving paragraphs and voxels with semantically similar explanations, e.g. directions and locations. Plot shows one subject (S02), similar plots for other subjects are shown in Section A.4. (c) After running the fMRI driving experiment, we found that a key factor determining whether a voxel is driven well by the GCT framework was the stability score for the voxel, i.e. the correlation between the n-gram rankings provided by the LLaMA-based encoding model and the OPT-based encoding model. (d) Another important factor for eliciting increased driving responses is the presence of key n-grams in the driving paragraphs. These n-grams induce a standard hemodynamic response curve that peaks at around 6 seconds, yielding a significant increase ($p=0.009$; one-sided t-test). (e) Finally, to test whether the driving results were sensitive to the particular voxels that were selected, we evaluated whether the GCT stories drove alternative voxels in each subject that were assigned the same explanation as the target voxels being driven. Both the targed voxels and alternative voxels also showed significantly increased driving responses ($p<0.05$; permutation test, FDR-corrected).
@@
The experiments thus far showed that GCT can effectively explain and drive a variety of locations throughout cortex, but fails to successfully drive some voxels. To understand these limitations, we evaluated factors that influence driving performance, beginning with factors surrounding the generation of story stimuli.
@@
Given that driving failures are neither the result of stimulus-explanation mismatches nor stimulus-model mismatches, we concluded that they must largely stem from limitations in the encoding models. We only applied GCT to voxels with high encoding model test performance, but this does not guarantee that the encoding model is accurate. Model performance is measured on a dataset that is limited in scope and that comprises real, noisy data. Both of these factors could cause a voxel to appear well-predicted despite having an inaccurate encoding model, or to appear poorly-predicted despite having an accurate encoding model. Since these limitations cannot be addressed directly, we investigated factors that improve confidence in the generated explanations by assessing their stability, a key principle underlying effective statistical interpretation according to the predictability, computability, and stability (PCS) framework [14, 34].
@@ 8   Discussion
This work is related to work that has generated stimuli to drive neurons in the visual cortex [43, 15, 44], as well as more recent work that has used decoding to provide insight into the selectivity of the visual cortex [45]. In language fMRI, the closest work is Tuckute et al. 2023, which demonstrated the ability to drive the language network as a whole (rather than individual voxels or ROIs) using sampling techniques as demonstrated in 13. In comparison to these prior works, ours is the first to demonstrate driving at an explanatory level, not only identifying stimuli that can drive a particular voxel or ROI, but also hypothesizing the shared semantic features that cause the underlying activity. This distinction provides an important scientific benefit, allowing us to automate the discovery of new and unexpected regions of interest for further study. As encoding models improve, the reliability of such discoveries will naturally increase in coincidence.
@@
Our work can be applied to generating a new class of stimuli for online, causal experiments of language processing. We consider this work a piece of a larger closed-loop system that continually refines its understanding of brain selectivity by generating hypotheses, testing these hypotheses with new experiments, and refining the hypotheses based on the experimental results. As an example, such a framework could be an especially effective tool in lesion case studies where general observations must be made efficiently from relatively little data.
@@
While effective, GCT has several limitations. Most notably, it focuses on a single explanation, missing voxel activity that is polysemantic, i.e. that is not driven well by any single explanation. See Section A.2 for a more detailed discussion of this limitation and how future work might overcome it. Second, the effectiveness of GCT depends on staying within the manifold of stimuli trained on during encoding model training. This limitation is explored further in A.6 where we show that poorly generated, off-manifold stimuli are more difficult to predict and yield inferior driving outcomes. GCT is also sensitive to the properties of the underlying encoding models–using different encoding models could yield different explanations which might be equally valid. Thus we must be cautious not to interpret the explanations given by GCT as uniquely capturing the functional properties of a voxel or ROI. However, this issue will become less severe over time as better and more diverse encoding models become available.
@@
The work here opens the door to a great deal of future research. One avenue should explore whether the degree of granularity of the explanation that GCT generates is appropriate for cortical language selectivity. For example, one of the explanations proposed by GCT is birthdays, but it would be somewhat surprising to have a specifically birthday-selective area in human cortex. Rather, it is more plausible that such a voxel is selective for something more generic, such as social gatherings [46]. Such ambiguities are of the type that are liable to be resolved by higher-fidelity modelling or iterative causal experiments, such as the hypothesis-refinement feedback system proposed above. GCT is also fundamentally limited by the spatiotemporal granularity of the scanning modality we test it on, fMRI. Applying GCT with faster or more fine-grained scanning modalities, such as 7T laminar fMRI, ECoG, or NeuroPixels, could reveal selectivity patterns that cannot be effectively observed through fMRI.
@@
GCT demonstrates a general framework for aligning the data-driven predictions of models to the theory-oriented hypotheses of classical science. In this paper, we show that applying techniques from modern LLM interpretability research can improve our understanding of cognitive brain processes. Through simple prompting-based methods, we successfully recover semantic selectivity hypotheses that have previously been the providence of painstaking controlled experiments created over decades. Moreover, we open the door to many more similar results including the identification of micro ROIs in poorly understood brain areas. Our framework is extensible to other modalities such as vision and - requiring only naturalistic data from a single subject - can be easily adapted for use in case study analyses. More generally, the high-level idea of GCT, using generative models as a tool to operationalize imprecise hypotheses about the natural world, has general applications well outside the bounds of neuroscience entirely. GCT will help move toward a future with increasingly nuanced, operationalized, and falsifiable explanations in our scientific discourse.
@@ 9   Methods
Framework for generating explanations GCT yields a short, natural-language explanation describing what elicits the strongest response from each of the 1,500 selected voxel encoding models. To obtain this explanation, we follow the summarize and score (SASC) framework introduced in previous work [17] (see Fig. 1b). SASC first generates candidate explanations (using GPT-4 [16]) based on the n-grams ($n=1,2,3$) that elicit the most positive response from the LLaMA encoding model. Each candidate explanation is then evaluated by generating synthetic data based on the explanation and testing the response of the encoding model to the data.
@@
Before running follow-up experiments, we checked that each paragraph’s text matched its generated explanation (Fig. 4a). This match was measured by prompting an LLM [49] to evaluate the fraction of trigrams in the paragraph that are relevant to the paragraph’s generating explanation. We also validated that each story paragraph drives the encoding model for its corresponding voxel (Fig. 4b). We generated 8 different stories by changing the random seed for each subject and kept the best 2 for S01, the best 6 for S02 (the pilot subject), and the best 2 for S03. During presentation, these newly presented stories were visually presented at approximately conversational cadence, instead of being listened to, as in the original encoding setting.
@@
Permutation testing details The main GCT driving results in the paper evaluate whether a particular voxel (or region of interest) yielded an increased average response to the driving paragraphs, i.e. the synthetic paragraphs generated specifically to drive that voxel. The permutation tests we conducted compared this average response to the voxel’s average response to other random paragraphs in the experiment. Beyond permutation testing, GCT can be viewed more generally as a causal test where the treatment variable is the explanation underlying the driving paragraph and the outcome is the average observed fMRI response, though it may not perfectly match the assumptions underlying common frameworks such as the Neyman-Rubin causal model [50].
@@ A.1 Prompting details
OPA Direction and location descriptions Avoid mentioning any specific location names (like “New York” or “Europe”).
OPA-only Self-reflection and growth Avoid mentioning any specific location names (like “New York” or “Europe”).
PPA Scenes and settings Avoid mentioning any specific location names (like “New York” or “Europe”).
PPA-only Unappetizing foods Avoid mentioning any specific location names (like “New York” or “Europe”).
@@ A.2 Driving polysemantic voxels
We additionally run experiments seeking to drive multiple descriptions corresponding to the same voxel. These experiments generate and test GCT stories in the same manner as testing individual explanations (Fig. 1c) but differ in the way they source voxel descriptions. In our first investigation (with subject S02), we select two explanations that both come from SASC, in the case that its summaries of the top ngrams yield two distinct explanations. In our second investigation (now with subject S03), we instead select two explanations that each come from a different encoding model. In both cases, each story tests 8 voxels with two explanations each. We average the results over two stories per subject.
@@ A.3 Story-level breakdowns
In this section, we show results for GCT at the level of individual stories. Fig. 6 shows the mean driving voxel response across all subjects and settings. Fig. 7 stratifies these synthetic stories based on the underlying prompt used to generate the stories and Fig. 8 shows results for these stories at driving voxel clusters rather than individual voxels.
@@
Table 2: Head motion for different subjects during GCT stories. S01 shows substantially larger motion than the other two subjects.
@@ A.5 Selectively driving pairs of voxels
Useful and explanatory scientific theories are robust to unrelated distributional shifts. For example, if a region or voxel is selective for food concepts, then that region should remain selective for food concepts even if those concepts cooccur with unrelated semantic concepts. In order to test this independence of our generated explanations, we test whether GCT can be used to simultaneously drive pairs of voxels in different areas across the brain.
@@ A.6 Checkerboard driving
A.7 Location-ROI extended results
Figure 14: Extended results for Fig. 2c for subject S03 in addition to S02. We focused on three ROIs that are known to have similar selectivity for place concepts: retrosplenial cortex (RSC), the parahippocampal place area (PPA), and the occipital place area (OPA). When explanations were generated for each ROI independently we found that each ROI was driven by all three driving paragraphs (left side). To distinguish these ROIs, we used GCT to find new explanations and construct stories that would selectively drive each area while suppressing the other two. Testing these stories in an fMRI experiment showed that in subject S02 we succeeded in finding selective explanations for two ROIs: RSC is selectively driven by location names and PPA by unappetizing foods. However, the explanation for OPA, spatial positioning & directions, drove responses in all three ROIs (right side). In subject S03, we succeed for all three ROIs. GCT explanations are the same for 4 of the 6 ROIs between the two subjects.
A.8 Language network driving extended results
Figure 15: Driving explanations and driving scores for selectively driving the five language network regions (i.e. driving the difference between a region and the average of the others) in Fig. 3c for S02 (left) and S03 (right). Diagonal elements show driving scores corresponding to the driving explanation and are generally not very positive.
Figure 16: Explanations and driving scores for the five language network regions in Fig. 3c for S02 and S03. Region responses show considerable similarity within each subject, but major differences across the two subjects.
A.9 Semantic explanation average contrast maps [^1] Full set of flatmaps can also be found at github.com/microsoft/automated-explanations